# Fixed-sparsity matrix approximation from matrix-vector products

### v1

@@ A.2 Anti-concentration for sums
Proposition 5 (Berry–Esseen; see e.g. [Ser80, §1.9]).
@@
We now argue the third absolute moments are bounded. Note that $x_{\ell}$ and $y_{\ell}$ are both normally distributed with mean zero and variance at most $1$ (since $\|\mathbf{u}\|_{2}\leq 1$ and $\|\mathbf{v}\|_{2}\leq 1$). Then $x_{\ell}y_{\ell}$ is sub-exponential with constant width parameter [Ver18, Lemma 2.7.7], and hence $\operatorname{\mathbb{E}}[|x_{\ell}y_{\ell}-\operatorname{\mathbb{E}}[x_{\ell}y_{\ell}]|^{3}]\leq\rho$ for some $\rho$.
@@ A.3 Other facts
For $d\geq 1$, suppose $\mathbf{G}\sim\operatorname{Gaussian}(d)$. Then $\operatorname{\mathbb{E}}[\|\mathbf{G}^{\mathsf{T}}\mathbf{G}\|_{\mathsf{F}}^{2}]\leq 3d^{3}$.
@@
Write $\mathbf{A}=\mathbf{G}^{\mathsf{T}}\mathbf{G}$. The diagonal entries of $\mathbf{A}$ are distributed as Chi-squared random variables with $d$ degrees of freedom. These entries have mean $d$ and variance $2d$. Likewise, the off-diagonal entries of $\mathbf{A}$ are distributed as the inner product of two independent standard normal Gaussian vectors of length $d$. These entries therefore have mean zero and variance $d$. Therefore, for $i\neq j$,
$\operatorname{\mathbb{E}}\bigl{[}[\mathbf{A}]_{i,i}^{2}\bigr{]}=\operatorname{\mathbb{V}}\bigl{[}[\mathbf{A}]_{i,i}\bigr{]}+\operatorname{\mathbb{E}}\bigl{[}[\mathbf{A}]_{i,i}\bigr{]}^{2}=2d+d^{2},\qquad\operatorname{\mathbb{E}}\bigl{[}[\mathbf{A}]_{i,j}^{2}\bigr{]}=\operatorname{\mathbb{V}}\bigl{[}[\mathbf{A}]_{i,j}\bigr{]}=d.$
By the linearity of expectation,
$\operatorname{\mathbb{E}}\bigl{[}\|\mathbf{A}\|_{\mathsf{F}}^{2}\bigr{]}=\sum_{i=1}^{d}\sum_{j=1}^{d}\operatorname{\mathbb{E}}\bigl{[}[\mathbf{A}]_{i,j}^{2}\bigr{]}=d\cdot(2d+d^{2})+(d^{2}-d)\cdot d\leq 3d^{3}.\qed$
Appendix B High probability algorithm
@@
A standard result [AV79, Prop 2.4 (a)] asserts that with $q=19/20$,

### v3

@@ B.2 Anti-concentration for sums
Proposition 5 (Berry–Esseen; see e.g. [serfling_80, §1.9]).
@@
We now argue the third absolute moments are bounded. Note that $x_{\ell}$ and $y_{\ell}$ are both normally distributed with mean zero and variance at most $1$ (since $\|\mathbf{u}\|_{2}\leq 1$ and $\|\mathbf{v}\|_{2}\leq 1$). Then $x_{\ell}y_{\ell}$ is sub-exponential with constant width parameter [vershynin_18, Lemma 2.7.7], and hence $\operatorname{\mathbb{E}}[|x_{\ell}y_{\ell}-\operatorname{\mathbb{E}}[x_{\ell}y_{\ell}]|^{3}]\leq\rho$ for some $\rho$.
@@ B.3 Other facts
For $r,d\geq 1$, suppose $\mathbf{G}\sim\operatorname{Gaussian}(r,d)$. Then
$\operatorname{\mathbb{E}}\big{[}\|\mathbf{I}\circ\mathbf{G}^{\mathsf{T}}\mathbf{G}\|_{\mathsf{F}}^{2}\big{]}=d(2r+r^{2}),\qquad\operatorname{\mathbb{E}}\big{[}\|\mathbf{G}^{\mathsf{T}}\mathbf{G}-\mathbf{I}\circ\mathbf{G}^{\mathsf{T}}\mathbf{G}\|_{\mathsf{F}}^{2}\big{]}=(d^{2}-d)r.$
@@
Write $\mathbf{A}=\mathbf{G}^{\mathsf{T}}\mathbf{G}$. The diagonal entries of $\mathbf{A}$ are distributed as Chi-squared random variables with $r$ degrees of freedom. These entries have mean $r$ and variance $2r$. Likewise, the off-diagonal entries of $\mathbf{A}$ are distributed as the inner product of two independent standard normal Gaussian vectors of length $r$. These entries therefore have mean zero and variance $r$. Therefore, for $i\neq j$,
$\operatorname{\mathbb{E}}\bigl{[}[\mathbf{A}]_{i,i}^{2}\bigr{]}=\operatorname{\mathbb{V}}\bigl{[}[\mathbf{A}]_{i,i}\bigr{]}+\operatorname{\mathbb{E}}\bigl{[}[\mathbf{A}]_{i,i}\bigr{]}^{2}=2r+r^{2},\qquad\operatorname{\mathbb{E}}\bigl{[}[\mathbf{A}]_{i,j}^{2}\bigr{]}=\operatorname{\mathbb{V}}\bigl{[}[\mathbf{A}]_{i,j}\bigr{]}=r.$
The result follows by linearity of expectation. ∎
Appendix C High probability algorithm
@@
A standard result [angluin_valiant_79, Prop 2.4 (a)] asserts that with $q=19/20$,